# Nellie: A Neuro-Symbolic Inference Engine for Grounded, Compositional, and Explainable Reasoning

### v4

@@ 5.3 Tree Error Analysis
We find that Nellie can produce high quality proofs of correct hypotheses; examples are shown in Figures 5 (left) and §F. However, the system can also generate proofs for incorrect answers that bypass its entailment filters. We can diagnose (and perhaps annotate) error patterns to address in future work. We list a few categories, depicted in Figure 5 (right). A common pattern of error is Hypernym Property Inheritance Errors, in which the model assumes that a member of a taxonomic class has a property of their hypernym, but the property is not universal. A colloquial example is inferring penguins can fly from penguins are birds & birds can fly. We also observe an amount of Errors from Semantic Drift (Khashabi et al., 2019) between inference hops, culminating in RTE model false positives. E.g., in block $2$ of Fig 5 (right), which object loses temperature during heat transfer changes between hops. Figure 7 shows the distribution of these errors on a set of 136 incorrect answers from EntailmentBank. The predominant error case is a failure to find any proof of the correct option.
@@ 6 Conclusion
We have proposed a reimagined version of a classical expert system that relies on the inferential power of LLMs rather than handcrafted rules. Nellie has the skeleton of a symbolic theorem prover, but the provided rules invoke neural predicates that allow for systematic reasoning over NL statements. In order to dynamically generate inferences, we introduce two mechanisms for knowledge-guided generation: conditioning decompositions on retrieval or model-selected inference templates. Our search algorithm undergirds Nellie, the first explainable reasoning system that performs end-to-end QA by searching for grounded entailment trees. We find that Nellie equals or exceeds the performance of comparable XQA systems, while providing explanations with the simultaneous guarantees of logical directionality and groundedness in human-provided text. This work thus suggests a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
@@ Appendix A Model Details and Training
We train the components of Nellie to be able to answer questions in the Science QA domain. The different neural modules are trained on reformulations of existing datasets for scientific reasoning. We use the following sources: WorldTree Explanations [Xie et al., 2020] and EntailmentBank [Dalvi et al., 2021] are used as described in the paper. eQASC [Jhamtani and Clark, 2020] and SciTaiL [Khot et al., 2018] are datasets of two- and one-premise entailment questions.
QA2D Our QA2D model is a T5-11B [Raffel et al., 2020] fine-tuned by [Tafjord et al., 2022] on the original QA2D dataset [Demszky et al., 2018].
Retrieval Module The dense retrieval model is a SentenceTransformers Siamese BERT-Network encoder [Reimers and Gurevych, 2019] pretrained on the MS MARCO IR corpus [Nguyen et al., 2016], which we fine-tune via ranking loss to maximize the cosine similarity between a hypothesis and its support facts. We gather $h,f_{1}$ from the WorldTree, EntailmentBank, and eQASC training sets, then sample random negatives during training.
Dynamic Rule Generator The DRG is a T5-3B model [Raffel et al., 2020] fine-tuned on $h\rightarrow[f_{1},f_{2}]$ pairs drawn from eQASC, which are numerous (26K) but noisy, and the binary composition steps from the EntailmentBank training set, which are few (3.8K) but ultimately rooted in facts from WorldTree. To make the DRG able to condition on templates, we adopt the span masking strategy from [Raffel et al., 2020] to create a randomly masked version $\hat{f}_{1}$ of each $f_{1}$ and then train the model on $h,\hat{f}_{1}\rightarrow f_{1},f_{2}$ pairs. For each original triplet, we create one training example with a template, and one in which $\hat{f}_{1}$ is an empty <mask>, giving the model the dual capacity for non-template-generation.
Entailment Filters We find that a mixture of strategies helps prioritize precision over recall when filtering candidates. For 1-premise filtering for unification and 2-premise filtering for rule generation, we fine-tune a separate pair of models consisting of a SentenceTransformer cross-encoder classifier and a T5 seq2seq. The former is trained via classification loss, the latter via sequence cross entropy loss. The 1-premise filters are trained on the SciTail dataset, the 2-premise filters on eQASC and EntailmentBank. For 2-premise entailment, we additionally add a “self-ask” filter, which following [Tafjord et al., 2022] is a T5 model fine-tuned to produce its 0-to-1 confidence that a statement is true. If the confidence is below 0.5 for either $f_{1}$ or $f_{2}$, we filter the pair. As in [Tafjord et al., 2022], we train our DRG seq2seq in a multitask fashion so that it serves as both the rule generator as well as the “self-ask” filter and (half of) the 2-premise entailment filter.
@@ Appendix B Hyperparameters
We used Optuna [Akiba et al., 2019] to find the number of facts for retrieval-conditioned generation, the $p$ value for nucleus sampling, the entailment threshold, and the number of candidates generated at each recursion step. We used as our objective the overall QA accuracy on the EntailmentBank dev set. We ran it for about 300 trials, and found that the most important parameters were the entailment threshold (too low = too many bad proofs, too high = no proofs) and the nucleus $p$ (less than 0.8 reduced QA performance by 10-15. Experiments were run on 24GB Quadro RTX 6000s.
@@ Appendix E Search Algorithm
Algorithm 1 depicts Nellie’s search algorithm, the logic of which is implemented in DeepProbLog [Manhaeve et al., 2018]. The search is parametrized by
  1. 1.
    A maximum number of proofs $m$ at which to cut off searching. We set this to $10$ for top-level queries and $2$ for recursive subqueries.
  2. 2.
    Candidate generation rates $n_{v}$ for vanilla nucleus-sampled decompositions, $n_{t}$ for template-conditioned decompositions, and $n_{r}$ for retrieval-conditioned generations. We set these to $200$, $250$ and $200$, respectively. [^11] Due to batching, these correspond to 1, 2, and 1 total calls to HuggingFace’s generate function. Upon removing exact match duplicates, a call to RuleGen produces about $400$ candidates.
  3. 3.
    The number of support facts $n_{f}$ to retrieve at each call to $\textsc{Retrieve}_{K}$, which we set to 15.
  4. 4.
    Entailment scoring module $s_{e}(\cdot)$, which is a separate NLI cross-encoder model for single- and double-premise entailments.
1:procedure Prove($H$, $m$) $\rightarrow$ proofs $tree^{(1)}(H),\dots tree^{(m)}(H)$ scores $s^{(1)}(H),\dots s^{(m)}{H}$:
@@
6:   ii. generate candidate decompositions: // If $H$ does not unify against $K$, decompose it
7:       decompositions $D={\color[rgb]{0,0,1}\textsc{RuleGen}}(H)$
8:   iii. check whether any decompositions pairs both unify: // this step amortizes premise lookups
9:       proofs = []; scores = []
10:       forall $(p_{1},p_{2})\in D$ do
11:               if $p_{1}\in K$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$ // if $p_{1}$ already grounded, its subtree is leaf
12:               then proofs += ($(\{p_{1},(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
13:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$ // tree score is min of scores used to construct it
14:               elif $\exists p_{1}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{1})$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$
15:               then proofs += ($(\{(p_{1}^{\prime}\vdash{}p_{1}),(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
16:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}^{\prime}\vdash{}p_{1}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
17:       if proofs not empty
18:       then return proofs, scores
19:   iv. recur on generated premises: // if still no proof, need to recur
20:       proofs = []; scores = []
21:       forall $(p_{1},p_{2})\in D$ do
22:               if ${\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H)>\max(\text{scores})$  and  $\text{len}(\text{proofs})<m$ // only recur if chance of better score
23:               then $(\text{proof}_{p_{1}},S_{p_{1}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{1},1)$
24:                   $(\text{proof}_{p_{2}},S_{p_{2}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{2},1)$
25:                   recursive score $S_{r}=\min(S_{p_{1}},S_{p_{2}},{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
26:                   if $S_{r}>\max(\text{scores})$
27:                   then proofs += ($(\{\text{proof}_{p_{1}},\text{proof}_{p_{2}}\}\vdash{}H)$
28:                      scores += $S_{r}$
29:       return proofs, scores
30:
31:procedure PremiseLookup(H) $\rightarrow$ premises $P$ s.t. $\forall p\in P,p\vdash{}H$ // try to ground H in K via entailment
32:       candidate premises $P={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$ // retrieve candidate premises s.t. $p\vdash{}H$
@@
34:               $P=f(P,H)$ // only keep premises that pass all filters
35:       return $P$
36:
37:procedure RuleGen(H) $\rightarrow$ decompositions $D=(p_{1}^{1},p_{2}^{1}),\dots(p_{1}^{c},p_{2}^{c})$ // find candidate pairs s.t. $p_{1}\land p_{2}\vdash{}H$
38:   D = []
39:   $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,n=n_{v})$
40:   WTTemplates = CaseBasedRetrieveTemplate(H)
41:   outputs per template $o_{t}=n_{t}/\text{len}({\color[rgb]{0.56,1,0.26}\textsc{WTTemplates}})$
42:   forall $t\in{\color[rgb]{0.56,1,0.26}\textsc{WTTemplates}}$ do // generate candidates for each of 50 templates in WTTemplates
43:       $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}([H\ t],n=o_{t})$ // T5 model accepts optional $t$ appended to $H$
44:   retrieved first premises $P_{1}={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$
45:   outputs per premise $o_{r}=n_{r}/\text{len}(P_{1})$
46:   forall $p_{K}\in P_{1}$ do // generate candidates with retrieved support facts as first premise
47:         $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,\text{prefix}=p_{K},n=o_{r})$ // Force decode $p_{K}$ then generate a $p_{2}$
@@
49:               $D=f(D,H)$ // only keep premise pairs that pass all filters
50:         return $D$
Algorithm 1 Nellie’s backchaining algorithm, coded in Prolog, for searching for proof list $tree^{(1)}(H),\dots tree^{(m)}(H)$ with scores $s^{(1)}(H),\dots s^{(m)}(H)$ for a hypothesis $H$ against factbase $K$. The rule generator produces $n_{v}$ candidates via vanilla nucleus sampling, $n_{t}$ candidates via sampling from templates, and $n_{r}$ candidates whose first fact is retrieved from K. All Generate, Retrieve, and entailment scorer $\textsc{Score}_{e}$ calls in forall loops are batched. All module outputs are cached. OnePremiseFilters and TwoPremiseFilters  are ordered lists of neural entailment models.
Appendix F Example Proofs
Figure 9 illustrates more proofs generated by Nellie for questions from EntailmentBank and WorldTree. Blue-outlined nodes are hypotheses generated by the QA2D model for the correct answer highted in green. Grey nodes are generated subqueries; tan nodes are leaves from Nellie’s knowledge corpus. The green node is a question setup text that grounds a contextual subquery via entailment.
Question
Nellie Proof
Cheetahs have come close to extinction due to hunting, drought, and disease. There is now very little genetic variation in cheetah populations. Which of the following is a result of the limited genetic variation in the current cheetah populations compared to earlier cheetah populations with more variation? (A) Cheetahs in current populations are more resistant to new diseases. (B) The survival rate of young cheetahs is increased in current populations. (C) Cheetahs in current populations are less able to interbreed with other species. (D) The current cheetah populations are less likely to be able to adapt to environmental changes.
In a city near the ocean, fog often forms on summer mornings. Which of the following statements best explains how this fog forms? (A) Ocean water evaporates and then condenses in the air. (B) Crashing waves spray tiny drops of ocean water into the air. (C) Water runoff moves toward the ocean and collects near the shore. (D) Rain clouds move in from the ocean and evaporate as they reach the shore.
The snowshoe hare sheds its fur twice a year. In the summer, the fur of the hare is brown. In the winter, the fur is white. Which of these statements best explains the advantage of shedding fur? (A) Shedding fur keeps the hare clean. (B) Shedding fur helps the hare move quickly. (C) Shedding fur keeps the hare’s home warm. (D) Shedding fur helps the hare blend into its habitat.
Scientists use large optical telescopes to obtain information about the planets in the solar system. What wavelengths of electromagnetic radiation provide this information? (A) gamma radiation (B) infrared radiation (C) radio waves (D) visible light
Figure 9: Example multiple-choice questions with NELLIE’s answer and corresponding proof.

### v5

@@ 5.3 Tree Error Analysis
We find that Nellie can produce high quality proofs of correct hypotheses; examples are shown in Figure 8 (left) and §F. However, the system can also generate proofs for incorrect answers that bypass its entailment filters. We can diagnose (and perhaps annotate) error patterns to address in future work. We list a few categories, depicted in Figure 8 (right). A common pattern of error is Hypernym Property Inheritance Errors (also known as the “fallacy of the undistributed middle”), in which the model assumes that a member of a taxonomic class has a property of their hypernym, but the property is not universal. A colloquial example is inferring penguins can fly from penguins are birds & birds can fly. We also observe an amount of Errors from Semantic Drift Khashabi et al. (2019) between inference hops, culminating in RTE model false positives. E.g., in block $2$ of Fig 8 (right), which object loses temperature during heat transfer changes between hops. Figure 9 shows the distribution of these errors on a set of 136 incorrect answers from EntailmentBank. The predominant error case is a failure to find any proof of the correct option.
Correct Answer Tree Analysis
To investigate whether Nellie reaches a right solution with right vs wrong reasoning, we manually inspected for reasoning errors in 50 correct answer trees produced by NELLIE. We found that 39/50 (78%) of trees were perfectly acceptable. Most of the 11 unacceptable trees were due to incompleteness at one decomposition step.
@@ 6 Conclusion
We propose a reimagined version of a classical expert system that relies on the inferential power of LLMs rather than handcrafted rules. Nellie has the skeleton of a symbolic theorem prover, but the provided rules invoke neural predicates that allow for systematic reasoning over NL statements. To dynamically generate inferences, we introduce two mechanisms for knowledge-guided generation: conditioning decompositions on retrieval or model-selected inference templates. Our search algorithm undergirds Nellie, an explainable reasoning system that performs end-to-end QA by searching for grounded entailment trees. We find that Nellie equals or exceeds the performance of comparable XQA systems, while providing explanations with the simultaneous guarantees of logical directionality and groundedness in human-provided text. This work thus suggests a new way to jointly reap the benefits of both modern neural methods and traditional reasoning.
Acknowledgements
Thanks to Nils Holzenberger, Kyle Richardson, Elias Stengel-Eskin, Marc Marone, Kate Sanders, Rachel Wicks, and Andrew Blair-Stanek for comments on earlier drafts. Thank you to Abbey Coogan for feedback on figure design.
@@ Appendix A Model Details and Training
We train the components of Nellie to be able to answer questions in the Science QA domain. The different neural modules are trained on reformulations of existing datasets for scientific reasoning. We use the following sources: WorldTree Explanations Xie et al. [2020] and EntailmentBank Dalvi et al. [2021] are used as described in the paper. eQASC Jhamtani and Clark [2020] and SciTaiL Khot et al. [2018] are datasets of two- and one-premise entailment questions.
QA2D Our QA2D model is a T5-11B Raffel et al. [2020] fine-tuned by Tafjord et al. [2022] on the original QA2D dataset Demszky et al. [2018].
Retrieval Module
The dense retrieval model is a SentenceTransformers Siamese BERT-Network encoder Reimers and Gurevych [2019] pretrained on the MS MARCO IR corpus Nguyen et al. [2016], which we fine-tune via ranking loss to maximize the cosine similarity between a hypothesis and its support facts. We gather $h,f_{1}$ from the WorldTree, EntailmentBank, and eQASC training sets, then sample random negatives during training.
Dynamic Rule Generator
The DRG is a T5-3B model Raffel et al. [2020] fine-tuned on $h\rightarrow[f_{1},f_{2}]$ pairs drawn from eQASC, which are numerous (26K) but noisy, and the binary composition steps from the EntailmentBank training set, which are few (3.8K) but ultimately rooted in facts from WorldTree. To make the DRG able to condition on templates, we adopt the span masking strategy from Raffel et al. [2020] to create a randomly masked version $\hat{f}_{1}$ of each $f_{1}$ and then train the model on $h,\hat{f}_{1}\rightarrow f_{1},f_{2}$ pairs. For each original triplet, we create one training example with a template, and one in which $\hat{f}_{1}$ is an empty <mask>, giving the model the dual capacity for non-template-generation.
Entailment Filters
We find that a mixture of strategies helps prioritize precision over recall when filtering candidates. For 1-premise filtering for unification and 2-premise filtering for rule generation, we fine-tune a separate pair of models consisting of a SentenceTransformer cross-encoder classifier and a T5 seq2seq. The former is trained via classification loss, the latter via sequence cross entropy loss. The 1-premise filters are trained on the SciTail dataset, the 2-premise filters on eQASC and EntailmentBank. For 2-premise entailment, we additionally add a “self-ask” filter, which following Tafjord et al. [2022] is a T5 model fine-tuned to produce its 0-to-1 confidence that a statement is true. If the confidence is below 0.5 for either $f_{1}$ or $f_{2}$, we filter the pair. As in Tafjord et al. [2022], we train our DRG seq2seq in a multitask fashion so that it serves as both the rule generator as well as the “self-ask” filter and (half of) the 2-premise entailment filter.
@@ Appendix B Hyperparameters
We used Optuna Akiba et al. [2019] to find the number of facts for retrieval-conditioned generation, the $p$ value for nucleus sampling, the entailment threshold, and the number of candidates generated at each recursion step. We used as our objective the overall QA accuracy on the EntailmentBank dev set. We ran it for about 300 trials, and found that the most important parameters were the entailment threshold (too low = too many bad proofs, too high = no proofs) and the nucleus $p$ (less than 0.8 reduced QA performance by 10-15. Experiments were run on 24GB Quadro RTX 6000s.
@@ Appendix E Search Algorithm
Algorithm 1 depicts Nellie’s search algorithm, the logic of which is implemented in a meta-interpreter coded in DeepProbLog Manhaeve et al. [2018].
Appendix F Example Proofs
Figure 11 illustrates more proofs generated by Nellie for questions from EntailmentBank and WorldTree.
Appendix G Impact of Design Choices
Nellie is comprised of numerous modules that are very difficult to evaluate in isolation given a lack of in-domain test sets for tasks like QA2D, decomposition generation ($H\rightarrow F1,F2$) compositional RTE ($H\Leftarrow F1,F2$), single-premise RTE ($H\Leftarrow F1$), and fact retrieval. Below, we enumerate some modeling choices, alternatives we tried, and the extent to which we observed them impact the system.
  • QA2D:
    Nellie uses:
    Tafjord et al. [2022]’s T5-11B model.
    We also tried: using GPT-J Wang and Komatsuzaki [2021] with in-context-learning exemplars drawn from EntailmentBank. This results in 10+ point drop in QA, mainly because (A) EntailmentBank hypotheses are designed to be generics instead of question-conditioned statements and (B) GPT-J tended to generate true statements even for wrong answer options.
  • Decomposition generation ($H\rightarrow F1,F2$):
    Nellie uses: A T5-3B model trained in a multi-angle fashion similar to Tafjord et al. [2022], using additional data for compositional entailment and decomposition generation.
    We also tried: Using a T5-Large model trained only on the decomposition data. This allowed us to generate far more decomposition candidates per hypothesis, but rate of obtaining a good decomposition was far lower. This led to 5+ point drop in QA.
  • Compositional RTE ($H\Leftarrow F1,F2$):
    Nellie uses:
    A sequence of (1) The same T5-3B model as for generation and (2) a cross-encoder trained on eQASC and EntailmentBank.
    We also tried: (A) using a bi-encoder trained via contrastive loss (this was not as effective at ruling out bad decompositions), (B) not training the bi-encoder to condition on the question, and (C) only training on EntailmentBank (as Tafjord et al. [2022] do). All three variants reduce QA by 3-7%.
  • Single-premise RTE ($H\Leftarrow F1$):
    Nellie uses:
    (1) A cross-encoder trained on Sci-Tail
    We also tried: an off-the-shelf RTE model [^11] https://huggingface.co/cross-encoder/nli-deberta-v3-base, which had too high a false positive rate to be effective.
  • Fact Retrieval:
    Nellie uses:
    a Sentence-Transformers bi-encoder pretrained on MS MARCO and fine-tuned on $(H,F1)$ pairs drawn from WorldTree, EntailmentBank, and eQASC.
    We also tried: the same bi-encoder off-the-shelf without finetuning. This only dropped performance by 3-4 points.
1:procedure Prove($H$, $m$, $d$) $\rightarrow$ proofs $tree^{(1)}(H),\dots tree^{(m)}(H)$ scores $s^{(1)}(H),\dots s^{(m^{\prime})}{(H)}$:
@@ Appendix G Impact of Design Choices
6:       elif $d=d_{\text{max}}$
7:       then Fail
8:   ii. generate candidate decompositions: // If $H$ does not unify against $K$, decompose it
9:       decompositions $D={\color[rgb]{0,0,1}\textsc{RuleGen}}(H)$
10:   iii. check whether any decompositions pairs both unify: // this step amortizes premise lookups
11:       proofs = []; scores = []
12:       forall $(p_{1},p_{2})\in D$ do
13:               if $p_{1}\in K$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$ // if $p_{1}$ already grounded, its subtree is leaf
14:               then proofs += ($(\{p_{1},(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
15:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$ // tree score is min of scores used to construct it
16:               elif $\exists p_{1}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{1})$  and  $\exists p_{2}^{\prime}\in{\color[rgb]{0,0,1}\textsc{PremiseLookup}}(p_{2})$
17:               then proofs += ($(\{(p_{1}^{\prime}\vdash{}p_{1}),(p_{2}^{\prime}\vdash{}p_{2})\}\vdash{}H)$
18:                   scores += $\min({\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}^{\prime}\vdash{}p_{1}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{2}^{\prime}\vdash{}p_{2}),{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
19:       if proofs not empty
20:       then return proofs, scores
21:   iv. recur on generated premises: // if still no proof, need to recur
22:       proofs = []; scores = []
23:       forall $(p_{1},p_{2})\in D$ do
24:               if ${\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H)>\max(\text{scores})$  and  $\text{len}(\text{proofs})<m$ // only recur if chance of better score
25:               then $(\text{proof}_{p_{1}},S_{p_{1}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{1},1,d+1)$
26:                   $(\text{proof}_{p_{2}},S_{p_{2}})={\color[rgb]{0,0,1}\textsc{Prove}}(p_{2},1,d+1)$
27:                   recursive score $S_{r}=\min(S_{p_{1}},S_{p_{2}},{\color[rgb]{1,.5,0}\textsc{Score}_{e}}(p_{1}\land p_{2}\vdash{}H))$
28:                   if $S_{r}>\max(\text{scores})$
29:                   then proofs += ($(\{\text{proof}_{p_{1}},\text{proof}_{p_{2}}\}\vdash{}H)$
30:                      scores += $S_{r}$
31:       return proofs, scores
32:
33:procedure PremiseLookup(H) $\rightarrow$ premises $P$ s.t. $\forall p\in P,p\vdash{}H$ // try to ground H in K via weak entailment
34:       candidate premises $P={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$ // retrieve candidate premises s.t. $p\vdash{}H$
@@
36:               $P=f(P,H)$ // only keep premises that pass all filters
37:       return $P$
38:
39:procedure RuleGen($H$) $\rightarrow$ decompositions $D=(p_{1}^{1},p_{2}^{1}),\dots(p_{1}^{c},p_{2}^{c})$ // find candidate pairs s.t. $p_{1}\land p_{2}\vdash{}H$
40:       $D$ = []
41:       $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,n=n_{v})$
42:       WTTemplates = CaseBasedRetrieveTemplate(H)
43:       outputs per template $o_{t}=n_{t}/\text{len}({\color[rgb]{0.51953125,0.69921875,0.421875}\textsc{WTTemplates}})$
44:       forall $t\in{\color[rgb]{0.51953125,0.69921875,0.421875}\textsc{WTTemplates}}$ do // generate candidates for each of 50 templates in WTTemplates
45:               $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}([H\ t],n=o_{t})$ // T5 model accepts optional $t$ appended to $H$
46:       retrieved first premises $P_{1}={\color[rgb]{1,.5,0}\textsc{Retrieve}_{K}}(H,n_{f})$
47:       outputs per premise $o_{r}=n_{r}/\text{len}(P_{1})$
48:       forall $p_{K}\in P_{1}$ do // generate candidates with retrieved support facts as first premise
49:               $D$ += ${\color[rgb]{1,.5,0}\textsc{Generate}}(H,\text{prefix}=p_{K},n=o_{r})$ // Force decode $p_{K}$ then generate a $p_{2}$
@@
51:               $D=f(D,H)$ // only keep premise pairs that pass all filters
52:       return $D$
[…]
Question Nellie Proof
In what way can forest fires affect the lithosphere? (A) increasing soil erosion (B) causing air pollution (C) increasing seismic activity (D) causing violent dust storms Refer to caption
The snowshoe hare sheds its fur twice a year. In the summer, the fur of the hare is brown. In the winter, the fur is white. Which of these statements best explains the advantage of shedding fur? (A) Shedding fur keeps the hare clean. (B) Shedding fur helps the hare move quickly. (C) Shedding fur keeps the hare’s home warm. (D) Shedding fur helps the hare blend into its habitat. Refer to caption
Scientists use large optical telescopes to obtain information about the planets in the solar system. What wavelengths of electromagnetic radiation provide this information? (A) gamma radiation (B) infrared radiation (C) radio waves (D) visible light Refer to caption
Figure 11: Example multiple-choice questions with NELLIE’s answer and corresponding proof. Blue-outlined nodes are hypotheses generated by the QA2D model for the correct answer highted in green. Grey nodes are generated subqueries; tan nodes are leaves from Nellie’s knowledge corpus.